OpenStudy (anonymous):

The perimeter of a retangler is 194 inches. The length exceeds the width by 29 inches. Find the width and the lenght please.

OpenStudy (anonymous):

Let say x is the breadth

OpenStudy (anonymous):

\[2(x+29)=194\]

OpenStudy (anonymous):

\[x+29=97\] \[x=97-29=>68\]

OpenStudy (anonymous):

So the length is \[68+29=97\]

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

ok whats the width sir

OpenStudy (anonymous):

Its the same as x i.e. 68

OpenStudy (anonymous):

I put 97 it said something.

OpenStudy (anonymous):

Ooops! I am very sorry

OpenStudy (anonymous):

Provided you the wrong equation at the first step

OpenStudy (anonymous):

\[2(x+x+29)=194\]

OpenStudy (anonymous):

Now solve for x thats all

OpenStudy (anonymous):

x would be width, and x+29 would be length

OpenStudy (anonymous):

Is it okay now?

OpenStudy (anonymous):

Bye now, got work to do..............